# SLRQA: A Sparse Low-Rank Quaternion Model for Color Image Processing with Convergence Analysis

### v1

@@ 5 An Algorithm for SLRQA-NF model
The augmented Lagrangian function of SLRQA-NF (4) is given by:
@@
~$\tilde{\mathcal{L}}_{\beta}(\mathbf{\dot{Z}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})=\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\left\langle\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{Z}}-\mathbf{\dot{W}}\right\rangle$ (53)
@@
$+\frac{\beta}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{W}}\|^{2}_{\mathrm{F}}+\left\langle\mathbf{\dot{\Lambda}}_{2},\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\right\rangle+\frac{\beta}{2}\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$
$=\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\lambda p(\mathbf{\dot{W}})+\frac{\beta}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{W}}+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{1}\|_{\mathrm{F}}^{2}$
$+\frac{\beta}{2}\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{2}\|_{\mathrm{F}}^{2}-\frac{1}{2\beta}\|\mathbf{\dot{\Lambda}}_{1}\|_{\mathrm{F}}^{2}-\frac{1}{2\beta}\|\mathbf{\dot{\Lambda}}_{2}\|^{2}_{\mathrm{F}},$
where $\beta>0$ is the penalty parameter, $\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2}\in\mathbb{H}^{m\times n}$ are the Lagrange multipliers.
In Algorithm 3, the subproblem of $\mathbf{\dot{Z}}$ in Step 5 is the proximal ADMM step i.e. the minimization of ~$\tilde{\mathcal{L}}_{\beta}(\mathbf{\dot{Z}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})$ for fixed $\mathbf{\dot{W}}_{k},\mathbf{\dot{\Lambda}}_{1,k},\mathbf{\dot{\Lambda}}_{2,k}$ plus a proximal term $\frac{L_{1,k}}{2}\|\mathbf{\dot{Z}}-\mathbf{\dot{Z}}_{k}\|_{\mathrm{F}}^{2}.$ Since the subproblem is equivalent to
@@
$\min_{\mathbf{\dot{Z}}}\,\sum_{i}\phi(\sigma_{i}(\mathbf{\dot{Z}}),\gamma)+\frac{\beta+L_{1,k}}{2}\left\|\mathbf{\dot{Z}}-\frac{1}{\beta+L_{1,k}}\left(\beta\mathbf{\dot{W}}_{k}+\mathbf{\dot{\Lambda}}_{1,k}+L_{1,k}\mathbf{\dot{Z}}_{k}\right)\right\|_{\mathrm{F}}^{2},$
Algorithm 2 can also be used to obtain $\mathbf{\dot{Z}}_{k+1}$.
In Step 6 of Algorithm 3, $\mathbf{\dot{W}}$, according to the linearization strategy in [21], a proximal gradient method is used in Step 6 to linearize the second order term. The penalty parameter of the second order term is denoted by $\beta\eta.$ To be specific, the quadratic term $\frac{\beta}{2}\|\mathbf{\dot{W}}-\mathbf{\dot{Z}}-\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{1}\|_{\mathrm{F}}^{2}+\frac{\beta}{2}\|\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}))-\mathcal{P}_{\Omega}(\mathbf{\dot{Y}})\|_{\mathrm{F}}^{2}$ in ~$\tilde{\mathcal{L}}_{\beta}(\mathbf{\dot{Z}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})$ is replaced by the linearized term plus a regularized term $\beta\left\langle\mathbf{\dot{W}}_{k}-\mathbf{\dot{Z}}_{k+1}-\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{1,k},\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle+\beta\left\langle\mathcal{W}\mathcal{A}^{\#}(\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{Y}})+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{2,k},\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\rangle$. Hence, the update of $\mathbf{\dot{W}}_{k+1}$ is given by
@@
$\mathbf{\dot{W}}_{k+1}=\mathbf{\dot{W}}_{k}-\frac{\mathcal{W}\mathcal{A}^{\#}(\mathcal{P}_{\Omega}(\mathcal{A}\mathcal{W}^{\#}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{Y}})+\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{2,k})+\mathbf{\dot{W}}_{k}-\mathbf{\dot{Z}}_{k+1}-\frac{1}{\beta}\mathbf{\dot{\Lambda}}_{1,k}}{\eta}-\frac{\lambda}{\beta\eta}\nabla p(\mathbf{\dot{W}}_{k}).$ (54)
Step 7 and 8 is used to update the Lagrange multipliers. Finally, the recovered image ${\mathbf{\dot{X}}}_{*}$ is given by ${\mathbf{\dot{X}}}_{*}=\mathcal{W}^{\#}(\mathbf{\dot{Z}}_{*})$.

### v2

@@ 4.2 Global Convergence for Algorithm 3
The global convergence is established for a generalization of SLRQA-NF, i.e. Model (11). The augmented Lagrangian function of (11) is
@@
$\mathcal{L}_{\beta_{1},\beta_{2}}({\mathbf{\dot{X}}},\mathbf{\dot{W}},\mathbf{\dot{\Lambda}}_{1},\mathbf{\dot{\Lambda}}_{2})=f({\mathbf{\dot{X}}})+g(\mathbf{\dot{W}})+\left\langle\mathbf{\dot{\Lambda}}_{1},\mathcal{C}_{1}({\mathbf{\dot{X}}})+\mathcal{B}_{1}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{1}\right\rangle+\left\langle\mathbf{\dot{\Lambda}}_{2},\mathcal{B}_{2}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{2}\right\rangle$
$\qquad\qquad\qquad\qquad\qquad\qquad\quad+\frac{\beta_{1}}{2}\|\mathcal{C}_{1}({\mathbf{\dot{X}}})+\mathcal{B}_{1}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{1}\|^{2}_{\mathrm{F}}+\frac{\beta_{2}}{2}\|\mathcal{B}_{2}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{2}\|_{\mathrm{F}}^{2}$
@@
$\qquad\qquad=f({\mathbf{\dot{X}}})+g(\mathbf{\dot{W}})+\frac{\beta_{1}}{2}\left\|\mathcal{C}_{1}({\mathbf{\dot{X}}})+\mathcal{B}_{1}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{1}+\frac{\mathbf{\dot{\Lambda}}_{1}}{\beta_{1}}\right\|_{\mathrm{F}}^{2}+\frac{\beta_{2}}{2}\left\|\mathcal{B}_{2}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{2}+\frac{\mathbf{\dot{\Lambda}}_{2}}{\beta_{2}}\right\|^{2}$
@@
$\qquad\qquad\qquad\quad-\frac{1}{2\beta_{1}}\left\|\mathbf{\dot{\Lambda}}_{1}\right\|_{\mathrm{F}}^{2}-\frac{1}{2\beta_{2}}\left\|\mathbf{\dot{\Lambda}}_{2}\right\|_{\mathrm{F}}^{2}.$
The iteration scheme for the proposed PL-ADMM-NF is
@@
${\mathbf{\dot{X}}}_{k+1}$ $\in\mathop{\mathrm{arg\,min}}_{{\mathbf{\dot{X}}}}f({\mathbf{\dot{X}}})+\frac{\beta_{1}}{2}\left\|\mathcal{C}_{1}({\mathbf{\dot{X}}})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k})-\mathbf{\dot{B}}_{1}+\frac{\mathbf{\dot{\Lambda}}_{k,1}}{\beta_{1}}\right\|_{\mathrm{F}}^{2}+\frac{L_{k,1}}{2}\|{\mathbf{\dot{X}}}-{\mathbf{\dot{X}}}_{k}\|_{\mathrm{F}}^{2},$ (59)
$\mathbf{\dot{W}}_{k+1}$ $\in\mathop{\mathrm{arg\,min}}_{\mathbf{\dot{W}}}\left\langle\nabla g(\mathbf{\dot{W}}_{k}),\mathbf{\dot{W}}\right\rangle+\frac{\beta_{1}}{2}\left\|\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{1}+\frac{\mathbf{\dot{\Lambda}}_{k,1}}{\beta_{1}}\right\|_{\mathrm{F}}^{2}$
$\qquad\qquad\qquad+\frac{\beta_{2}}{2}\|\mathcal{B}_{2}(\mathbf{\dot{W}})-\mathbf{\dot{B}}_{2}+\frac{\mathbf{\dot{\Lambda}}_{k,2}}{\beta_{2}}\|_{\mathrm{F}}^{2}+\frac{L_{k,2}}{2}\left\|\mathbf{\dot{W}}-\mathbf{\dot{W}}_{k}\right\|^{2}_{\mathrm{F}},$
$\mathbf{\dot{\Lambda}}_{k+1,1}$ $=\mathbf{\dot{\Lambda}}_{k,1}+\mu\beta_{1}(\mathcal{C}_{1}({\mathbf{\dot{X}}}_{k+1})+\mathcal{B}_{1}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{1}),$
$\mathbf{\dot{\Lambda}}_{k+1,2}$ $=\mathbf{\dot{\Lambda}}_{k,2}+\mu\beta_{2}(\mathcal{B}_{2}(\mathbf{\dot{W}}_{k+1})-\mathbf{\dot{B}}_{2}).$
The following assumptions are made for the convergence of PL-ADMM-NF.
Assumption 4.2.
  1. B1
    $f({\mathbf{\dot{X}}})$ and $g(\mathbf{\dot{W}})$ are bounded from below and have $L_{f}$ and $L_{g}$ Lipschitz continuous gradients respectively, i.e., for every ${\mathbf{\dot{X}}}_{1},{\mathbf{\dot{X}}}_{2},\mathbf{\dot{W}}_{1}$, and $\mathbf{\dot{W}}_{2}$, it holds that
    $\|\nabla g(\mathbf{\dot{W}}_{1})-\nabla g(\mathbf{\dot{W}}_{2})\|_{\mathrm{F}}\leq L_{g}\|\mathbf{\dot{W}}_{1}-\mathbf{\dot{W}}_{2}\|_{\mathrm{F}},$ (60)
    $\|\nabla f({\mathbf{\dot{X}}}_{1})-\nabla f({\mathbf{\dot{X}}}_{2})\|_{\mathrm{F}}\leq L_{f}\|{\mathbf{\dot{X}}}_{1}-{\mathbf{\dot{X}}}_{2}\|_{\mathrm{F}}.$
  2. B2
    The matrix $\mathcal{C}_{1}^{\#}\mathcal{C}_{1}$ is full rank or $f({\mathbf{\dot{X}}})$ is coercive, $\mathcal{B}_{2}^{\#}\mathcal{B}_{2}$ is full rank, ${\rm range}(\mathcal{B}_{1})\subseteq{\rm range}(\mathcal{C}_{1})$, $\mathbf{\dot{B}}_{1}\in{\rm range}(\mathcal{C}_{1})$, and $\mathbf{\dot{B}}_{2}\in{\rm range}(\mathcal{B}_{2})$.
  3. B3
    The parameters $\beta_{1},\beta_{2}>0$, $\mu\in(0,2)$, and there exist three constants $a_{1}>0,a_{2}>0$, and $r>1$ such that
    $(q_{1}^{-}-3r\theta_{1})\mathbf{I}\succeq a_{1}\mathbf{I}\,\quad\mbox{and}\,\quad q_{2}^{-}\mathbf{I}+\beta_{2}\mathcal{B}_{2}^{\#}\mathcal{B}_{2}+\beta_{1}\mathcal{B}_{1}^{\#}\mathcal{B}_{1}-(3r\theta_{2}+L_{g})\mathbf{I}\succeq a_{2}\mathbf{I},$ (61)
    where $\mathbf{I}$ is the identity matrix with correct size and $\theta_{1}=\theta_{1,1}+\theta_{2,1},\theta_{2}=\theta_{1,2}+\theta_{2,2}$,
    $\theta_{1,1}=\frac{4(q_{1}+L_{f})^{2}\mu}{\rho(\mu)^{2}\beta_{1}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},\quad\theta_{1,2}=\frac{4\beta_{1}\|\mathcal{B}_{1}\|^{2}\|\mathcal{C}_{1}\|^{2}}{\rho(\mu)\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},$ (62)
    $\theta_{2,1}=\frac{3\mu\|\mathcal{B}_{1}\|^{2}(4q_{1}^{2}+|1-\mu|^{2}(4(q_{1}+L_{f})^{2}))}{\beta_{2}\rho(\mu)^{4}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}},$
    $\theta_{2,2}=\frac{6(L_{g}+q_{2})^{2}\mu}{\beta_{2}\rho(\mu)^{2}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}}+\frac{3\mu\|\mathcal{B}_{1}\|^{2}(|1-\mu|^{2}+1)(4\beta_{1}^{2}\|\mathcal{B}_{1}\|^{2}\|\mathcal{C}_{1}\|^{2})}{\beta_{2}\rho(\mu)^{4}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}}.$
  4. B4
    The parameters $\beta_{1},\beta_{2}$ satisfies $\frac{\kappa}{2L_{f}}>\theta_{3,0}+\theta_{4,0},\frac{1}{2L_{g}}>\theta_{4,0},$ where $\kappa\in(0,1)$ is a given constant and
    $\theta_{3,0}=\frac{3}{2\beta_{1}\rho(\mu)\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}},\quad\theta_{4,0}=\max\left\{\frac{3}{2\beta_{2}\rho(\mu)\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}},\frac{3\|\mathcal{B}_{1}\|^{2}}{2\beta_{2}\rho(\mu)^{2}\lambda_{+}^{\mathcal{B}_{2}^{\#}\mathcal{B}_{2}}\lambda_{+}^{\mathcal{C}_{1}^{\#}\mathcal{C}_{1}}}\right\}.$ (63)
@@
For (B1), the differentiability of $f({\mathbf{\dot{X}}})$ and Lipschitz continuity of $\nabla f({\mathbf{\dot{X}}})$ in SLRQA-NF (4) comes from the differentiability of $\phi$ and Lipschitz differentiability of $\nabla\phi$. For more discussions on the Lipschitz continuity for spectral function, we refer the readers to [23, Corollary 2.5] and [13, Theorem 3.3].